OpenStudy (anonymous):

how would I show 7776 as a power using the base of 6?? I know it is 6^5 because I divided by 6 until I got 1, but is there a different way?

OpenStudy (anonymous):

you can cheat if you like

OpenStudy (anonymous):

with a calculator that is.

OpenStudy (anonymous):

if you didn't see immediately that \[6^5=2776\] you could solve \[6^x=2776\] for x by taking out your calculator and computing \[\log(2776)\div \log(6)\] and you will get 5 try it

OpenStudy (anonymous):

all typos there . should be \[6^x=7776\] \[x=\log(7776)\div \log(6)\]